# Guided Transfer Learning for Discrete Diffusion Models

### v1

@@ 4 Guided Transfer Learning for DDMs
$=\prod_{\ell=1}^{L}\frac{\Big[r_{\phi}\!\big(\mathbf{z}_{t,s^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\displaystyle\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}\Big[r_{\phi}\!\big(\mathbf{z}_{t,\tilde{s}^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}.$ (8)
A naive integration of this guidance into the sampling process incurs a computational complexity of $\mathcal{O}(|\mathcal{V}|\cdot L)$ per denoising step, since we must evaluate $r_{\phi}$ for all sequences obtained by replacing every possible token in $\mathcal{V}$ at each position $L$ in the denominator. Prior work (schiff2025simpleguidancemechanismsdiscrete; nisonoff2025unlockingguidancediscretestatespace) targets settings with modest vocabulary sizes and sequence lengths, where this cost is manageable. In our experiments with the Markov chain (Section 5.1) it is also not prohibitive. But for text applications, the cost becomes significant. (nisonoff2025unlockingguidancediscretestatespace) proposes a first-order Taylor approximation that reduces evaluations, but it is accurate primarily in the high-noise regime and degrades as $t$ decreases.
4.1 Scalability via Planner Sampling
In this section, we tackle the main bottleneck: integrating guidance efficiently at inference time. The gains from a set of simple techniques are summarized in Table 1.
@@ 4.1 Scalability via Planner Sampling
Ancestral Sampling
In the naive implementation, at all denoising steps, we compute logits from the denoiser and from the guidance network with $\mathcal{O}(|\mathcal{V}|\cdot L)$ complexity. Following (sahoo2024simpleeffectivemaskeddiffusion), in masked diffusion, we can remove explicit time conditioning because the number of masked tokens in the sequence already implies the timestep. This is not possible for uniform noise, which mixes token identities. We therefore cache logits across steps. If a step does not change the sequence, we reuse the logits from the previous step. The number of sampling steps where the network is called then becomes upper-bounded by the sequence length, since in the worst case, every position unmasks in a different step.
In addition, we introduce a top-$n$ style pruning mechanism for the ratio network. Specifically, we assume that (i) the denoiser distribution $p_{\theta}(\cdot\mid z_{t})$ is sufficiently concentrated, so that the true target distribution $q$ differs from $p$ only within the top-$n_{\text{ratio}}$ candidates, and (ii) the density ratio $r(z)=q(z)/p(z)$ is bounded, in the sense that there exists a constant $C<\infty$ such that for all tokens within the top-$n_{\text{ratio}}$ set of $p_{\theta}$, we have $\frac{q(z)}{p(z)}\leq C$. Outside this set, where $p(z)$ is already very small, no such bound is required. Consequently, we evaluate the ratio term only for the top $n_{\text{ratio}}$ candidates proposed by the denoiser at each position. This results in position-specific candidate sets in the denominator of equation (8) and reduces the computational cost of ratio evaluations to $\mathcal{O}(n_{\text{ratio}}\cdot L)$ with $n_{\text{ratio}}\leq|\mathcal{V}|$. In the experiments, the combination of top-$n_{\text{ratio}}$ selection and the guidance weight $\gamma$ can misallocate probability mass, which may cause the diffusion process to collapse into a fully masked sequence at the final step. Thus, a key assumption is that the probability of remaining a masked token must be equal between the target and source distributions. As a result, we enforce a position-independent normalization which implies the renormalized update over non-mask tokens, thus stabilizing the denoising diffusion process. Further details of this stabilizer are presented in the supplementary materials.
@@
We also apply our method to text in Section 5.2, where $n_{\text{ratio}}\ll\lvert\mathcal{V}\rvert$, yet sampling remains slow and costly mainly because of the $L$ in the complexity. To address this, we propose a planner-based algorithm. The planner $\rho_{\vartheta}$ is trained to predict the next denoising position in the masked sequence. Concretely, we replace the Gillespie process with $\tau$-leaping by a planner schedule that unmasks exactly one position per step, thereby fixing the number of denoising steps to $L$. In many implementations, $T\approx L$, so enforcing a hard budget of $L$ steps does not materially change the step count, while reducing guidance evaluations and wall-clock time. Pseudocode is provided in Algorithm 1. An important achievement here is that the loop in lines 4 to 7 can now be parallelized by vectorization on a single GPU. As a result, only one forward pass of the three networks is required.
For training the planner, we adopt a simplified training algorithm following (peng2025pathplanningmaskeddiffusion; liu2025thinkgeneratediscretediffusion). For ratio training, we use a discrete data training loss function building on the work of (ouyang2024transferlearningdiffusionmodels). Further details of the loss function and training algorithm are provided in the supplementary materials.
@@ 20 Refer to caption Refer to caption Refer to caption
We construct two Markov chains with distinct transition matrices. Let the probability state vector be $\mathbf{\pi}_{i}\in\Delta^{N-1}$ over a vocabulary of size $N=5$. The transition matrix of the target domain is $\mathbf{T}\in\mathbb{R}^{N\times N}$ and that of the source domain is $\mathbf{S}\in\mathbb{R}^{N\times N}$. We sample sequences of length $20$ from the target domain using the update rule $\mathbf{\pi}_{i}=\mathbf{T}\,\mathbf{\pi}_{i-1}$ with the initial state $\mathbf{\pi}_{0}$ set to the uniform vector. Sampling from the source domain proceeds in the same way.
Implementation
We use the sampling process of (sahoo2024simpleeffectivemaskeddiffusion) for the source, finetuned, and target discrete diffusion models. The backbone is a Diffusion Transformer (DiT) (peebles2023scalablediffusionmodelstransformers) with about five million learnable parameters. The ratio estimator and the classifier use about half as many parameters with a final mean and fully connected layer that maps the output to a scalar. The exact implementation appears in our codebase. For guided sampling, we follow the ancestral sampling method. Since the vocabulary is small we set $n_{\text{ratio}}=N$ and use $20$ sampling steps. We train a masked diffusion model on the source domain for $30$ epochs. We then compare a vanilla masked diffusion model trained directly on the target samples (60 epochs), a finetuned diffusion model (90 epochs), and our GTL method (60 epochs). The finetuned model and our method share the same source diffusion model.
@@ 5.2 Language Modeling
In this section, we apply our method to text data. We use the arXiv abstracts dataset (arxiv_org_submitters_2024), which contains 1.7 million arXiv articles with titles, authors, categories, and abstracts. We focus on Computer Science, Mathematics, and Physics. For data preparation, we concatenate all text and split it into segments of 512 tokens using the bert-base-uncased tokenizer with vocabulary size N=30,522N=30{,}522. This results in 285,946285{,}946 samples in Computer Science, 176,831176{,}831 in Mathematics, and 79,63179{,}631 in Physics. For each category we randomly select 10% of the samples for validation. We ensure that the three category subsets are disjoint. We define the union of Computer Science and Mathematics as the source domain and Physics as the target domain.
@@
We train the source diffusion model for 100,000 gradient steps using the DiT architecture (peebles2023scalablediffusionmodelstransformers) with 59.8M trainable parameters. A log–linear noise scheduler is used for both training and sampling. The finetuned model on the target domain is trained for 10,000 gradient steps. For our guided method, we initialize a ratio model with 4.1M parameters and train it for 5,000 gradient steps. Additional training details appear in the supplementary materials.
For each method we generate 128 sequences of length 512. We evaluate quality with MAUVE (pillutla2021mauvemeasuringgapneural), which captures the tradeoff between perplexity and diversity. Although generative perplexity (Gen. PPL) is often reported, we observed, consistent with prior work (wang2025remaskingdiscretediffusionmodels), that Gen. PPL can remain low even when the text has reduced utility, such as repeated words. In those cases, MAUVE is more sensitive and provides a clearer measure of quality. We also evaluate the domain of each sampled sequence using an independently trained classifier. Scores near zero indicate the target domain and scores near one indicate the source domain. The classifier follows the setup used for the ratio model in the supplementary materials with label smoothing set to $0.05$.
Behavior of $\gamma$ and $n_{\text{ratio}}$
@@
Method Comparison and Sensitivity to Data Fraction
@@ 6 Conclusion
Limitations
Overall, the proposed method is promising and applicable to a wide range of tasks involving domain shifts. However, it requires access to both source and target domains. Future work could explore conditional transfer for prompt-based or attribute-controlled generation. It would also be interesting to evaluate the learned ratio on quality-filtered tasks, such as reducing toxicity in language modeling.
7 Supplementary Materials for Guided Transfer Learning for Discrete Diffusion Models
Appendix A Proof of Theorem 1
@@
To conclude Lemma 2, we leverage the fact that the source and target domain share the same forward kernel in Equation 14 to 15 $q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=\frac{q(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})\cdot q(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})}{q(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$.
@@
For 27 to 28 we show that $\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}$:
@@
For 28 to 29 we show that $\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}=\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\bigl[q(\mathbf{x}_{0})/p(\mathbf{x}_{0})\bigr]$:
@@
Appendix B Guided $\tau$-Sampling
@@
Appendix C Training Details
@@
Appendix D Application to continuous-time Score-based Models
D.1 Background
(campbell2022continuoustimeframeworkdiscrete) introduces a continuous-time framework for discrete denoising models formulated as a continuous-time Markov chain (CTMC). Unlike the discrete-time setting of (austin2023structureddenoisingdiffusionmodels) with steps $t=0,1,\dots,T$, here $t\in[0,T]$ is continuous.

### v2

@@ 3 Guided Transfer Learning for DDMs
$q_{\theta,\phi}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)}\right)=\prod_{\ell=1}^{L}\frac{\Big[r_{\phi}\!\big(\mathbf{z}_{t,s^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\displaystyle\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}\Big[r_{\phi}\!\big(\mathbf{z}_{t,\tilde{s}^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}.$ (7)
A naive implementation of equation 7 is prohibitively expensive. For each masked position $\ell$, the denominator requires evaluating $r_{\phi}(\mathbf{z}_{t,\tilde{s}^{\ell}}^{(1:L)})$ for all $\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}$, resulting in $\mathcal{O}(L|\mathcal{V}|)$ ratio evaluations per denoising step. This is manageable for small $L$ or $|\mathcal{V}|$ (e.g., our Markov-chain setting), but becomes prohibitive for text.
@@
3.1 Scalability via Planner Sampling
Our goal is to reduce ratio-network evaluations from $\mathcal{O}(L|\mathcal{V}|)$ to a budget that is independent of $|\mathcal{V}|$ and ideally weakly dependent on $L$. Table 1 summarizes how caching, top-$n_{\text{ratio}}$, and planner sampling progressively reduce this cost.
Ancestral Sampling
In the naive implementation, at all denoising steps, we compute logits from the denoiser and from the guidance network with $\mathcal{O}(|\mathcal{V}|\cdot L)$ complexity.
Caching. In masked diffusion, the timestep is implicitly determined by the number of masked tokens, allowing us to remove explicit time conditioning from the networks (Sahoo et al., 2024). As a result, if a denoising step does not modify the sequence, the corresponding logits can be reused. By caching logits across steps, the number of network evaluations becomes upper-bounded by the number of unmasking events. In the worst case, each position unmasks once, yielding at most $L$ effective network calls. This optimization does not apply to uniform-noise diffusion, where token identities are mixed. Top-$n_{\text{ratio}}$ pruning. Although caching reduces the number of steps, each step still requires evaluating the ratio network over the full vocabulary. To address this, we restrict ratio evaluations to a small candidate set. We assume that the denoiser distribution $p_{\theta}(\cdot\mid\mathbf{z}_{t})$ is sufficiently concentrated, so that meaningful deviations between the source and target distributions occur primarily among the top-$n_{\text{ratio}}$ tokens. Formally, we assume that for these candidates the density ratio $r(z)=q(z)/p(z)$ is bounded by a constant $C<\infty$; no such bound is required outside this set, where $p(z)$ is already negligible. Under this assumption, we evaluate the ratio term only for the top-$n_{\text{ratio}}$ tokens proposed by the denoiser at each position. This produces position-specific candidate sets in the denominator of equation 7 and reduces the cost of ratio evaluations to $\mathcal{O}(n_{\text{ratio}}\,L)$, with $n_{\text{ratio}}\ll|\mathcal{V}|$ in practice.
Mask-probability stabilizer. The interaction between top-$n_{\text{ratio}}$ pruning and the guidance weight $\gamma$ can distort probability mass, occasionally causing the diffusion process to collapse to fully masked sequences. To prevent this failure mode, we enforce a position-independent constraint that preserves the probability of remaining masked across source and target distributions. This induces a renormalized update over non-mask tokens and stabilizes the denoising dynamics. Details of this stabilization are provided in the appendix.
@@ 3.1 Scalability via Planner Sampling
Even with top-$n_{\text{ratio}}$ pruning, the overall complexity still scales linearly with $L$. To further reduce cost, we introduce a planner-based sampler that selects a single position to update at each step. The planner $\rho_{\vartheta}$ is trained to predict the next denoising position in the masked sequence, replacing the stochastic Gillespie process with a deterministic $\tau$-leaping schedule. This strategy fixes the number of denoising steps to exactly $L$, which is comparable to standard masked diffusion where typically $T\approx L$, while substantially reducing guidance evaluations and wall-clock time. Algorithm 1 summarizes the resulting ratio-guided denoising step. Crucially, once the planner selects a position, ratio evaluations over the top-$n_{\text{ratio}}$ candidates can be fully vectorized. As a result, each denoising step requires only a single forward pass through the denoiser, ratio network, and planner. For planner training, we follow simplified strategies from Peng et al. (2025); Liu et al. (2025). The ratio network is trained using a discrete-data loss adapted from Ouyang et al. (2024). Additional training details are provided in the supplementary materials.
@@ 20 Refer to caption Refer to caption Refer to caption
We construct two Markov chains with distinct transition matrices. Let the probability state vector be $\mathbf{\pi}_{i}\in\Delta^{N-1}$ over a vocabulary of size $N=5$. The transition matrix of the target domain is $\mathbf{T}\in\mathbb{R}^{N\times N}$ and that of the source domain is $\mathbf{S}\in\mathbb{R}^{N\times N}$. We sample sequences of length $20$ from the target domain using the update rule $\mathbf{\pi}_{i}=\mathbf{T}\,\mathbf{\pi}_{i-1}$ with the initial state $\mathbf{\pi}_{0}$ set to the uniform vector. Sampling from the source domain proceeds in the same way. Implementation details can be found in the appendix E.1.
@@ 4.2 Language Modeling
In this section, we apply our method to text data. We use the arXiv abstracts dataset (arXiv.org submitters, 2024), which contains 1.7 million arXiv articles with titles, authors, categories, and abstracts. We focus on Computer Science, Mathematics, and Physics. For data preparation, we concatenate all text and split it into segments of 512 tokens using the bert-base-uncased tokenizer with vocabulary size N=30,522N=30{,}522. This results in 285,946285{,}946 samples in Computer Science, 176,831176{,}831 in Mathematics, and 79,63179{,}631 in Physics. For each category we randomly select 10% of the samples for validation. We ensure that the three category subsets are disjoint. We define the union of Computer Science and Mathematics as the source domain and Physics as the target domain.
@@
We train the source diffusion model for 100,000 gradient steps using the DiT architecture (Peebles and Xie, 2023) with 59.8M trainable parameters. A log–linear noise scheduler is used for both training and sampling. The finetuned model on the target domain is trained for 10,000 gradient steps. For our guided method, we initialize a ratio model with 4.1M parameters and train it for 5,000 gradient steps. Additional training details appear in the supplementary materials.
For each method we generate 128 sequences of length 512. We evaluate quality with MAUVE (Pillutla et al., 2021), which captures the tradeoff between perplexity and diversity. Although generative perplexity (Gen. PPL) is often reported, we observed, consistent with prior work (Wang et al., 2025), that Gen. PPL can remain low even when the text has reduced utility, such as repeated words. In those cases, MAUVE is more sensitive and provides a clearer measure of quality. We also evaluate the domain of each sampled sequence using an independently trained classifier. Scores near zero indicate the target domain and scores near one indicate the source domain. The classifier follows the setup used for the ratio model in the supplementary materials with label smoothing set to $0.05$.
4.2.1 Behavior of $\gamma$ and $n_{\text{ratio}}$
@@
4.2.2 Method Comparison and Sensitivity to Data Fraction
@@ 5 Conclusion
Limitations Overall, the proposed method is promising and applicable to a wide range of tasks involving domain shifts. However, it requires access to both source and target domains. Future work could explore conditional transfer for prompt-based or attribute-controlled generation. It would also be interesting to evaluate the learned ratio on quality-filtered tasks, such as reducing toxicity in language modeling.
Appendix A Appendix
Appendix B Related Work
Discrete Diffusion Models
Discrete diffusion models (DDMs) have become a standard choice for generative modeling over categorical sequences, including biological sequences (DNA, RNA, proteins) and text (Sarkar et al., 2024; Yu et al., 2025). Compared to approaches that add continuous noise to embeddings or latents, fully discrete formulations avoid the projection step back to tokens, which can introduce an embedding-to-token mismatch (Li et al., 2022b; Lovelace et al., 2023). Two families of discrete diffusion methods are especially common. The first is discrete-time diffusion in the D3PM framework, where a Markov corruption process is defined by transition matrices and a denoiser is trained to invert the forward kernels; simplified masked variants such as MLDM and MD4 fall into this category (Sahoo et al., 2024; Shi et al., 2025; Austin et al., 2023). The second is continuous-time discrete diffusion, which parameterizes a discrete score (a log-ratio between neighboring-time marginals) and induces a continuous-time Markov chain (CTMC) for generation (Lou et al., 2024; Meng et al., 2023; Campbell et al., 2022).
Guided Diffusion Models
Guidance is a widely used mechanism for steering diffusion sampling in both continuous and discrete domains. In discrete diffusion, classifier-free guidance typically combines conditional and unconditional predictions at inference time, requiring two forward passes of a single model trained on a mixture of conditional and unconditional examples (Huang and Tang, 2025). Classifier-based guidance is closer to our setting: it introduces an auxiliary classifier $p(y\mid x_{t},t)$ and reweights the reverse transition so that samples are biased toward the desired attribute y, while the denoiser itself is trained on the full corpus (Schiff et al., 2025; Nisonoff et al., 2025). A direct implementation scores token substitutions across all positions and vocabulary items, requiring $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per step, which is infeasible for long sequences and large vocabularies. Because each evaluation corresponds to a distinct single-token intervention on the full sequence, this cost is not removed by straightforward parallelization. To reduce this cost, (Nisonoff et al., 2025) proposes a first-order Taylor approximation that replaces many evaluations with a forward and backward pass. Related alternatives adjust sampling using inference-time optimization: TESS2 (Tae et al., 2025) performs gradient ascent on a reward with respect to the model logits, and energy-based diffusion language models define an energy over predicted clean sequences and use it to reweight denoising updates to reduce train–sample mismatch (Xu et al., 2025).
Transfer Learning
Fine-tuning and transfer learning share the goal of adapting a model to a new task or domain. Fine-tuning is computationally expensive because it updates the pretrained denoiser’s weights on new data. Prior work on continuous diffusion models explores reinforcement learning (Fan et al., 2023; Clark et al., 2024) and adapter-based fine-tuning (Moon et al., 2022; Xie et al., 2023a). In contrast, transfer learning for diffusion remains comparatively underexplored. Transfer Learning for Diffusion Models (TLDM) (Ouyang et al. (2024)) proposes a ratio-estimation technique for continuous score-based models,but it relies on continuous scores and is evaluated in low-dimensional settings. Our work closes the gap for discrete diffusion by deriving ratio-guided reverse transitions in both discrete-time, continuous-time and score based formulations for high dimansion application.
Appendix C Proof of Theorem 1
@@
To conclude Lemma 2, we leverage the fact that the source and target domain share the same forward kernel in Equation 13 to 14 $q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=\frac{q(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})\cdot q(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})}{q(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$.
@@
For 26 to 27 we show that $\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}$:
@@
For 27 to 28 we show that $\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}=\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\bigl[q(\mathbf{x}_{0})/p(\mathbf{x}_{0})\bigr]$:
@@
Appendix D Guided $\tau$-Sampling
@@
Appendix E Training Details
@@
E.1 Implementation on Markov Chain
We use the sampling process of (Sahoo et al., 2024) for the source, finetuned, and target discrete diffusion models. The backbone is a Diffusion Transformer (DiT) (Peebles and Xie, 2023) with about five million learnable parameters. The ratio estimator and the classifier use about half as many parameters with a final mean and fully connected layer that maps the output to a scalar. The exact implementation appears in our codebase. For guided sampling, we follow the ancestral sampling method. Since the vocabulary is small we set $n_{\text{ratio}}=N$ and use $20$ sampling steps. We train a masked diffusion model on the source domain for $30$ epochs. We then compare a vanilla masked diffusion model trained directly on the target samples (60 epochs), a finetuned diffusion model (90 epochs), and our GTL method (60 epochs). The finetuned model and our method share the same source diffusion model.
Appendix F Application to continuous-time Score-based Models
F.1 Background
Campbell et al. (2022) introduces a continuous-time framework for discrete denoising models formulated as a continuous-time Markov chain (CTMC). Unlike the discrete-time setting of (Austin et al., 2023) with steps $t=0,1,\dots,T$, here $t\in[0,T]$ is continuous.